A car is accelerated uniformly along the horizontal plane from 0 to 90km/hr during 5 seconds. Find at what angle to the horizon the force acting on the driver from the seat is directed during such acceleration 


1
Expert's answer
2021-03-14T19:13:21-0400

The force acting on the driver from the seat is the resultant force of two forces: gravitational and inertial. This force is directed with the angle to the horizon, that equals:


"\alpha=arctan(\frac{g}{a}),"


where g - the acceleration of gravitational force;

a - the acceleration of inertial force.


"a=\frac{v}{t}=\frac{90}{3.6\cdot5}=5\space m/s^2."


"\alpha=arctan(\frac{9.8}{5})=63\degree"


Answer: 63"\degree"